The diameter of an electric cable, Say x is assumed to be Continuous random a Variable with Pdf f(x) = 6x * (1 - x), 0 <= x <= 1 check that f(x) is a P.d.f. ,Determine a&b such that P(x<b) = P(x>b) ,Find P(0.1<x <0.3)

1
Expert's answer
2021-11-21T13:34:12-0500

a)


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{1}6x(1-x)dx"

"=[3x^2-2x^3]\begin{matrix}
 1 \\
 0
\end{matrix}=1"

b)


"P(X<b)=\displaystyle\int_{0}^{b}6x(1-x)dx"

"=[3x^2-2x^3]\begin{matrix}
 b \\
 0
\end{matrix}=3b^2-2b^3, 0<b<1"

"P(x>b)=1-P(X\leq b)"

"=1-(3b^2-2b^3)"

"P(x<b)=P(x>b)"

"3b^2-2b^3=1-(3b^2-2b^3)"

"4b^3-6b^2+1=0"

"4b^3-2b^2-(4b^2-1)=0"

"2b^2(2b-1)-(2b-1)(2b+1)=0"

"(2b-1)(2b^2-2b-1)=0"

"2b-1=0=>b_1=1\/2"

"2b^2-2b-1=0"

"D=(-2)^2-4(2)(-1)=12"

"b=\dfrac{2\pm\sqrt{12}}{2(2)}=\dfrac{1\pm\sqrt{3}}{2}"

"b_2=\dfrac{1-\sqrt{3}}{2}<0, b_3=\dfrac{1+\sqrt{3}}{2}>1"

Therefore "b=1\/2"


c)


"P(0.1<X<0.3)=\displaystyle\int_{0.1}^{0.3}6x(1-x)dx"

"=[3x^2-2x^3]\begin{matrix}
 0.3 \\
 0.1
\end{matrix}"

"=3(0.3)^2-2(0.3)^3-(3(0.1)^2-2(0.1)^3)"

"=0.188"